# Intrinsic Dimensionality as a Model-Free Measure of Class Imbalance

### v1

@@ A.1 Literature on Long-Tailed Visual Recognition
Other Approaches with alternative perspectives are also present: E.g., residual fusion of specialized heads [25], exploration of an optimal combination of existing techniques [20, 51], a specialized activation function [26] or distribution-aware expert learning [49, 54, 55, 27]. Moreover, model calibration techniques to overcome over-confidence show promising improvements [46, 22]. Complementarily, a branch of works focuses on learning unbiased feature representations, with supervised contrastive learning [57, 44, 50, 48] or data augmentation [43].
@@
A.2 Literature on Intrinsic Dimension Estimation
@@
Appendix B Pseudocode of Our Approach
@@
Appendix C Implementation, Training and Experiment Details
C.1 Dataset Details
@@
C.2 Implementation and Training Details
@@
C.3 Experiment Details
For our experiments using the Bag of Tricks [20] framework, we use the training configurations reported in Tab. 9. For experiments using Focal Loss [42], we additionally sample $\gamma\in\{1.0,2.0\}$.
@@ C.3 Experiment Details
Tab. 10 shows the training configuration for BCL [44] + ID on ImageNet-LT. BCL is trained with a base learning rate of 0.1 and a cosine annealing learning rate scheduler. Following [44], we use $\lambda=1.0$ and $\mu=0.35$ as the logit compensation and balanced contrastive learning loss weights respectively.
The training configurations of SURE [51] on the CIFAR-LT datasets are listed in Tab. 11. SURE is trained with sharpness-aware minimization [76] together with stochastic weight averaging [77] (SWA) for increased robustness. SWA is initialized at 120 epochs with a SWA-specific learning rate of 0.05. Mix-up augmentation is used with the mixup parameter sampled from $\mathrm{Beta}(\beta,\beta)$.
Tab. 12 shows the training configuration for GLMC [43] on ImageNet-LT and CIFAR-LT datasets. On CIFAR-LT datasets, the model is trained with the SGD optimizer with a momentum of 0.9 with a cosine annealing learning rate scheduler. Mix-up augmentation is used with the mixing parameter sampled from $\mathrm{Beta}(\beta,\beta)$.
@@
Appendix D Additional Results
@@
D.1 Rarity-based Analysis on One-Stage Model-Free Approaches
We further analyze the results of ID-based imbalance mitigation with one-stage model-free methods in a rarity-based grouping and present the results in Tab. 14. We sort the 100 classes of CIFAR-100 by decreasing example counts and group the 33/34/33 class partitions as head, middle, and tail classes respectively. ID-based reweighting and resampling both outperform alternative methods, and in particular the results show that ID-based mitigation primarily increases the performance of trained in the middle and tail classes of a dataset.
@@ D.1 Rarity-based Analysis on One-Stage Model-Free Approaches
D.2 Additional Analysis on FisherS ID Estimation
@@
D.3 Additional Ablation Results on the Choice of the ID Estimator
Along with the results reported in Section 6.4, we additionally perform an ablation for the choice of ID estimator with the ImageNet-LT dataset in Tab. 15, where the results indicate that the FisherS estimator performs better than alternative estimators when paired up with the BCL [44] method.
In general, our results indicate that there is no singular optimal choice of ID estimator for all cases, and the optimal estimator may be different for a particular dataset and a particular method.
Dataset: ImageNet-LT
BCL [44] 56.1
BCL + ID (w/ MLE) 54.4
BCL + ID (w/ TLE) 55.5
BCL + ID (w/ FisherS) 56.5
D.4 Additional Robustness Analysis of Our Method
@@
D.5 Failure Cases with Our Method
We also briefly explore cases where our ID-based method may fail to be effective in imbalance mitigation. Some examples of failure cases for our approach is when ID fails to be informative for a particular class; one way to assure this is to reverse or shuffle the ID values for each class in the dataset. Tab. 16 shows the result of mitigation with uninformative ID values on CIFAR-10, where the results trail behind standard CE-based training.
Dataset: CIFAR-10 CIFAR-100
CE 71.60 38.90
CSCE [8] 74.69 41.08
ID 76.44 41.65
ID (Reversed) 71.55 39.33
ID (Shuffled) 71.52 $\pm$ 0.85 38.84 $\pm$ 0.61
Table 16: Top-1 accuracy results with uninformative ID configurations for CIFAR datasets with an imbalance ratio of 100.
Figure 11: ID estimates with low sample sizes on CIFAR-10.

### v2

@@ B.1 Literature on Long-Tailed Visual Recognition
Other Approaches with alternative perspectives are also present: E.g., residual fusion of specialized heads [18], exploration of an optimal combination of existing techniques [76, 42], a specialized activation function [3] or distribution-aware expert learning [77, 69, 9, 74]. Moreover, model calibration techniques to overcome over-confidence show promising improvements [78, 73]. Complementarily, a branch of works focuses on learning unbiased feature representations, with supervised contrastive learning [19, 81, 20, 41] or data augmentation [24]. There also exist other approaches on visual recognition, such as those based on clustering and class prototypes of feature spaces [68, 43, 46].
@@
B.2 Literature on Intrinsic Dimension Estimation
@@
Appendix C Pseudocode of Our Approach
@@
Appendix D Implementation, Training and Experiment Details
D.1 Dataset Details
@@
D.2 Implementation and Training Details
@@
D.3 Experiment Details
For our experiments using the Bag of Tricks [76] framework, we use the training configurations reported in Table 12. For experiments using Focal Loss [44], we additionally sample $\gamma\in\{1.0,2.0\}$.
@@ D.3 Experiment Details
Table 13 shows the training configuration for BCL [81] + ID on ImageNet-LT. BCL is trained with a base learning rate of 0.1 and a cosine annealing learning rate scheduler. Following [81], we use $\lambda=1.0$ and $\mu=0.35$ as the logit compensation and balanced contrastive learning loss weights respectively.
The training configurations of SURE [42] on the CIFAR-LT datasets are listed in Table 14. SURE is trained with sharpness-aware minimization [28] together with stochastic weight averaging [35] (SWA) for increased robustness. SWA is initialized at 120 epochs with a SWA-specific learning rate of 0.05. Mix-up augmentation is used with the mixup parameter sampled from $\mathrm{Beta}(\beta,\beta)$.
Table 15 shows the training configuration for GLMC [24] on ImageNet-LT and CIFAR-LT datasets. On CIFAR-LT datasets, the model is trained with the SGD optimizer with a momentum of 0.9 with a cosine annealing learning rate scheduler. Mix-up augmentation is used with the mixing parameter sampled from $\mathrm{Beta}(\beta,\beta)$.
@@
D.4 Computational Overhead of ID Estimation
Here, we briefly illustrate the computational overhead of training with our method. Our method’s computational costs only include the calculation of ID: it does not have any additional costs for all mitigation methods (re-sampling, re-weighting, or re-margining). Table 16 includes a wall-clock based timer for estimating ID on various datasets using the FisherS [2] estimator, including an estimate of the overhead of ID estimation to total training. Here, our reference training schedule is a two-stage re-sampling approach for each respective dataset, and the overhead corresponds to the ratio of total ID estimation time to training a single model. We reiterate that ID only needs to be computed once per dataset, rather than being re-computed each training session; thus the computational cost of ID estimation is constant with respect to the length of the training stage.
Dataset Total Time Time Per Class Overhead
CIFAR-10-LT ($\rho=100$) 56.1 $\pm$ 1.6s 5.6s 0.05
CIFAR-100-LT ($\rho=100$) 10.7s $\pm$ 0.3s 0.1s 0.01
ImageNet-LT 11m2s $\pm$ 7s 0.6s 0.01
Places-LT 42m59s $\pm$ 33s 7.1s 0.03
Table 16: Compute table for ID estimation for various datasets.
Appendix E Additional Results
@@
E.1 Rarity-based Analysis on One-Stage Model-Free Approaches
We further analyze the results of ID-based imbalance mitigation with one-stage model-free methods in a rarity-based grouping and present the results in Table 18. We sort the 100 classes of CIFAR-100 by decreasing example counts and group the 33/34/33 class partitions as head, middle, and tail classes respectively. ID-based reweighting and resampling both outperform alternative methods, and in particular the results show that ID-based mitigation primarily increases the performance of trained in the middle and tail classes of a dataset.
@@ E.1 Rarity-based Analysis on One-Stage Model-Free Approaches
E.2 Additional Analysis on FisherS ID Estimation
@@
E.3 Additional Robustness Analysis of Our Method
@@
E.4 An Inspection of ID-Cardinality Fusion
In this section, we briefly consider a fusion of ID-based and cardinality-based weighting. Specifically, we introduce a re-weighting strategy with class-based weights $\tilde{w_{c}}$ defined as the multiplicative combination of FisherS-based ID weighting and class-based weighting: $\tilde{w_{c}}=N_{c}^{-\beta}d_{c}^{\gamma}$, with scaling parameters $\beta,\gamma\in[0,1]$. We then train ResNet-32 models with re-weighting on the CIFAR-10-LT dataset with these fusion of weights. Fig. 14 shows the results of these trained models as a heatmap of Top-1 accuracies. In particular, we observe that with these fusion of weights, increasing the importance of the ID-based weighting with parameter $\gamma$ has a larger positive impact on classification accuracy compared to the cardinality-based weighting with parameter $\beta$.
Figure 14: Heatmap of Top-1 accuracies for methods ID-cardinality fusion on CIFAR-10-LT with an imbalance ratio of $\rho=100$.